OpenStudy (anonymous):

x^2+y^2-2x-10y+2=0 find the circle's center and radius? is the center 1; -5 and the radius is square root 28? just wanted to make sure if i got the correct answer or not

OpenStudy (anonymous):

center is (1, 5) and radius is sqrt(24) = 2sqrt(6)

OpenStudy (anonymous):

can u explain how did u get sqrt 24

OpenStudy (anonymous):

\[x^2+y^2-2x-10y+2=0\] Start off by completing the square: \[\left(x^2-2x\right)+\left(y^2-10y\right)+2=0\\ \left(x^2-2x\color{red}{+1}\right)+\left(y^2-10y\color{blue}{+25}\right)+2\color{red}{-1}\color{blue}{-25}=0\\ (x-1)^2+(y-5)^2=24\\(x-1)^2+(y-5)^2=(\sqrt{24})^2\] Now compare to the general equation of the circle centered at \((h,k)\) with radius \(r\): \[(x-h)^2+(y-k)^2=r^2\]